# RobEthiChor: Automated Context-aware Ethics-based Negotiation for Autonomous Robots

### v1

@@ 6.1 ROS-based architecture
The Ethics Manager Node subscribes to the
/current_context and /ethic_profile topics to realize the selection of the active profile. When selected (as described in Section 5), the active profile is published on the /active_profile topic channel to make it available for negotiation.
@@ 6.2 Negotiation algorithm implementation
8:  send $quit$
9:  return  $no\_ag$
@@
[…]
7:   return  $lose$
@@
[…]
@@
[…]
@@ 7 Evaluation
In the following, we detail the experiments. The code, the experimentation settings, and the results are available in the provided replication package [^10] https://github.com/RoboChor/robethichor-ethics-based-negotiation.
@@ 7.1 EQ1 & EQ2: Negotiation scenarios
To realize the negotiation scenarios for EQ1 and EQ2, we defined a set of 12 dispositions that concern the tasks involved in the resource contention scenarios, plus 27 possible user status conditions that activate the defined dispositions. Starting from this common ground, we created 10 different personas, each representing a potential user, by eliciting their status and ethical preferences. Then, we built (i) the context models for the two negotiation contexts (i.e., the airport and the hospital) and, for each persona, (ii) an ethical profile that considers possibly varying dispositions for different contexts (by aptly assigning context-related grades to the dispositions, as in Table 1), and (iii) a user status (by setting to true the values associated with their conditions). By enumerating all the possible combinations of pairs of users (45 pairs), we manually built a ground truth by associating the expected negotiation outcome with each pair of users for each of the two negotiation contexts, according to their profiles and statuses. Then, we ran the negotiation for each of the pairs and repeated the runs four times per each context (45 possible pairs, 180 runs per context). The logs obtained after running the simulations were analyzed and compared with the ground truth to assess whether the system’s behavior reflects the user’s expectations and to measure the time required for running the negotiation.
@@
The results obtained through the automated negotiation implemented in RobEthiChor fully comply with the ones expected with the ground truth. The negotiation always showed the same outcome in each run that featured the same pair of users in a given context, regardless of their order in the pair, the robot acting on their behalf, and their starting roles (i.e., if a user started the negotiation as a sender or receiver). That is, for each pair in the airport context, and for each pair in the hospital context, either the same user always won the negotiation or no agreement was always reached.
@@
Figure 5 summarizes the results of the negotiation between the 45 possible combinations of personas: in the airport context (Figure 5(a)) an agreement was reached in 140 out of 180 runs (77.8%), while in 40 runs (22.2%), no agreement was reached, as expected from the ground truth data. It is worth remarking that no agreement was reached in 40 runs because, according to the defined personas and the derived ethical profiles, none of the two negotiating users (i.e., robots on their behalf) was willing to concede to the other the contended resource. The negotiation in the hospital context (Figure 5(b)) gave different results: here, the agreement was reached in 132 out the 180 runs (73.3%), while no agreement was reached in 48 runs (26.7%), as expected from the ground truth for this context. Moreover, we observed that, for 16 pairs of users, the negotiation results were different in the two contexts (i.e., a different agreement was reached). These different results highlight how the outcome of the negotiation between the same pair of users may be influenced by the context where the negotiation happens since, as explained in Sections 3 and 4, the grades in a given user’s ethical profile for different contexts may differ.
@@ 7.2 EQ3: Scalability experiments
[…]
@@
It is worth discussing that, in general, the settings where the negotiation time introduced a high overhead, which can be arguably considered unacceptable (e.g., more than 3 seconds) are arguably unlikely in a real-world scenario. As explained above, we ran these settings to stress the system beyond reasonable human-provided input sizes with the primary aim of assessing the scalability dimension. In fact, the aforementioned settings feature huge profiles that consider more than 100 dispositions and, more importantly, an extremely wide and detailed set of possible user conditions (100 or more), most of which hold (i.e., most of the user status conditions are set to true) in the current negotiation context (75% or more). From a practical point of view, this means that we are admitting the existence of users who activated, at the same time, most of the (hundreds) possible status conditions, even though some of them may be conflicting, e.g., child and elderly. As a reference, to make it realistic, personas described in Section 7.1 have, at most, 8 active conditions out of the 27 defined ($\sim$30of the total). Overall, among the 1500 negotiation runs in the on-robot deployment, 76.5% of them ended in less than 3 seconds; this value raises to 84.73% in the off-robot deployment over the same configurations (among the full set of 2500 negotiation runs in the off-robot, 62.3% of them ended within 3 seconds). However, the results discussed above hint at possible strategies to reduce the negotiation time even in the “extreme” settings where the negotiation overhead becomes unacceptable. Due to the nature of the overhead, which is strongly dependent on the network communication, the off-robot deployment allows reducing the negotiation time if compared to the on-robot deployment (see Table 2). In our experimental setting, the overhead is approximately halved when the negotiation is run locally on an external computer. Moreover, reducing the size of the offer set or the number of exchanged offers would be effective in further reducing the overhead. This can be achieved either by (i) disclosing more than one status condition in each offer or negotiation round, e.g., increasing the value of $i$ in Algorithm 1 and 1 by more than 1 each round, or (ii) introducing a time deadline, as further discussed in Section 9.2.
@@ 8 Discussion
In this section, we discuss relevant aspects of the proposed approach.

### v2

@@ 6.1 ROS-based architecture
The Ethics Manager Node subscribes to the /current_context and /ethic_profile topics to realize the selection of the active profile. When selected (as described in Section 5), the active profile is published on the /active_profile topic channel to make it available for negotiation.
@@ 6.2 Negotiation algorithm implementation
8:  send quitquit
9:  return no_agno\_ag
@@
6:   reply(acceptaccept)
7:   return loselose
@@
10: reply(rejectreject)
11:return sender_algorithm($\mathcal{O}$, $\bar{o}$, $i+1$, $o^{\star}$)
@@
The receiver subroutine in Algorithm 2 is executed by providing as input the set of offers $\mathcal{O}$, the offer $\overline{o}$, and the index of the negotiation step $i$. When the robot starts the negotiation as a receiver, the algorithm is called with i=0i=0888The index $i$ keeps track of the number of sent offers, since, when playing the sender role, robots send the offer $o_{i}$. The subroutine implements the decision-making algorithm according to case 1 in Formula (2). After receiving a non-empty offer $o^{\star}$, the robot running the receiver algorithm computes the utility value $\mathcal{U}$. If the utility value is greater than 0, the robot accepts the negotiation by replying with an accept message on the /negotiation topic, and the algorithm ends by returning loser (lines 4-9). Otherwise, the robot rejects the offer (line 10) and switches its role to be a sender in the next negotiation round (line 11).
@@ 7 Evaluation
In the following, we detail the experiments. The code, the experimentation settings, and the results are available in the provided replication package [^10] https://doi.org/10.5281/zenodo.17449923.
@@ 7.1 EQ1 & EQ2: Negotiation scenarios
To realize the negotiation scenarios for EQ1 and EQ2, we defined a set of 17 dispositions that concern the tasks involved in the resource contention scenarios, plus 27 possible user status conditions that activate the defined dispositions. Starting from this common ground, we created 10 different personas, each representing a potential user, by eliciting their status and ethical preferences. Then, we built (i) the context models for the two negotiation contexts (i.e., the airport and the hospital) and, for each persona, (ii) an ethical profile that considers possibly varying dispositions for different contexts (by aptly assigning context-related grades to the dispositions, as in Table 1), and (iii) a user status (by setting to true the values associated with their conditions). Figure 5 reports two of the defined personas (Alice and Bob). We provided a card for each persona summarizing the persona’s characterizing traits, a description of the needs as a system user, and the conditions (i.e., the user status) in the two negotiating contexts. The figure also reports the ethical profiles built for the two personas, showing the grades assigned to each of the dispositions in the airport (A) and hospital (H) contexts. By enumerating all the possible combinations of pairs of users (45 pairs), we manually built a ground truth by associating the expected negotiation outcome with each pair of users for each of the two negotiation contexts, according to their profiles and statuses. Then, we ran the negotiation for each of the pairs and repeated the runs four times per each context (45 possible pairs, 180 runs per context). The logs obtained after running the simulations were analyzed and compared with the ground truth to assess whether the system’s behavior reflects the user’s expectations and to measure the time required for running the negotiation.
Refer to caption ALICE Age: 29
Occupation: Professional Athlete
Personality: Disciplined • Independent • Resilient
Hobbies: Hiking • Yoga • Reading
Tech use: Comfortable with assistive tools, especially when practical
Traits: Alice is a 29-year-old professional athlete. She’s disciplined, active, and enjoys hiking, yoga, and reading in her downtime. Normally independent and confident, Alice prefers to manage things on her own. After a recent leg injury and learning she is pregnant, she became more aware of situations that require extra care and attention. She believes that people with visible health conditions, disabilities, or medical needs, and unaccompanied minors, should be given support first. She also understands how overwhelming less obvious issues like anxiety or crowd stress can be. Needs (Airport): For a medical check-up related to her pregnancy and recovery, Alice travels to another city. At the airport, she feels the strain of walking long distances through crowded areas. The noise and lack of clear signs make the experience more tiring than expected. Spotting a nearby assistant robot offering mobility help, she decides to use it, which makes moving easy for her and saves her energy. Needs (Hospital): Once in the city, Alice visits the hospital for her check-up. While she tries to stay calm and focused, the emotional weight of her physical recovery and pregnancy catches up with her. She feels vulnerable and more easily overwhelmed than usual. Despite this, she values when procedures are explained clearly and handled efficiently. Though she’s still not used to asking for help, moments like these have shown her the value of supportive tools that make the journey easier without compromising her independence.   Airport user status: Injured, Pregnant, Pre-natal care, Crowd Stress Hospital user status: Injured, Pregnant, Pre-natal care, Emotional distress, Crowd stress
Refer to caption BOB Age: 67
Occupation: Retired (former school teacher)
Personality: Calm • Patient • Respectful • Quietly observant
Hobbies: Gardening • Reading newspapers • Listening to classical music
Tech use: Basic use of mobile phone and tablet
Traits: Bob is 67 years old and recently retired. He enjoys gardening, reading, and quiet music. Though generally active, he has some age-related health issues and gets tired walking long distances. Recently, he also suffered a mild fall that left him with some lingering leg pain, making mobility more difficult than usual. English is not his first language, which sometimes makes travel challenging. He believes that people with health issues, the elderly, and non-native speakers should be given priority. While usually calm, he appreciates support when situations become confusing or rushed. Needs (Airport): Bob is traveling to another city for a medical check-up. At the airport, he feels a bit lost and pressed for time. Walking far is tiring, and the signs aren’t always clear. When he notices an assistant robot offering help, he decides to use it. It guides him to his gate smoothly and lowers his stress. Needs (Hospital): When Bob reaches the hospital, he stays patient and attentive. However, when the check-in process becomes unexpectedly long and the waiting area grows crowded, he starts to feel anxious and emotionally worn down. He believes that patients with medical needs or physical impairments should go first and appreciates clear communication from the staff. Though he doesn’t often ask for help, simple, supportive systems make a big difference in easing his discomfort.   Airport user status: Elderly, Age-related health issues, Rushing to flight, Non-native speaker Hospital user status: Elderly, Injured, Non-native speaker, Emotional distress, Age-related health issues
Disposition Description (Give precedence to …..) Alice’s Grades (A) Alice’s Grades (H) Bob’s Grades (A) Bob’s Grades (H) $d_{1}$ the injured 2 4 1 3 $d_{2}$ the pregnant 4 3 3 3 $d_{3}$ the elderly 3 3 4 1 $d_{4}$ the children 2 1 1 1 $d_{5}$ an accompanied minor 5 2 1 2 $d_{6}$ one with general medical needs 4 4 4 4 $d_{7}$ the non-native speaker 1 2 3 2 $d_{8}$ one to avoid non-obvious/ non-medical situations (anxiety, crowd stress, etc.) 3 3 1 1 $d_{9}$ one traveling with a pet 0 0 0 0 $d_{10}$ whom the general assistance is required (lost something, flight delay, luggage transfer, fragile luggage, is lost, etc.) 2 0 3 0 $d_{11}$ ones with time-sensitive emergencies (tight flight schedule, delayed arrival, unseen emergency). 2 0 3 0 $d_{12}$ the disabled (mobility issues, hearing or visual impairments, or other disabilities) 4 4 3 3 $d_{13}$ one accompanied by a dependent (child, elderly, or disabled person) 2 1 3 2 $d_{14}$ one experiencing emotional distress 0 3 0 3 $d_{15}$ one with infectious symptoms (e.g., fever, coughing) 0 4 0 4 $d_{16}$ one needing urgent diagnostic access 0 4 0 4 $d_{17}$ one recovering from surgery (post-op) 0 3 0 2
Figure 5: Overview of Alice and Bob personas.
@@
The results obtained through the automated negotiation implemented in RobEthiChor fully comply with the ones expected with the ground truth and align with the ethical considerations in the defined personas. The negotiation always showed the same outcome in each run that featured the same pair of users in a given context, regardless of their order in the pair, the robot acting on their behalf, and their starting roles (i.e., if a user started the negotiation as a sender or receiver). That is, for each pair in the airport context, and for each pair in the hospital context, either the same user always won the negotiation or no agreement was always reached.
@@
Figure 6 summarizes the results of the negotiation between the 45 possible combinations of personas: in the airport context (Figure 6(a)) an agreement was reached in 140 out of 180 runs (77.8%), while in 40 runs (22.2%), no agreement was reached, as expected from the ground truth data. It is worth remarking that no agreement was reached in 40 runs because, according to the defined personas and the derived ethical profiles, none of the two negotiating users (i.e., robots on their behalf) was willing to concede to the other the contended resource. The negotiation in the hospital context (Figure 6(b)) gave different results: here, the agreement was reached in 132 out the 180 runs (73.3%), while no agreement was reached in 48 runs (26.7%), as expected from the ground truth for this context. Moreover, we observed that, for 16 pairs of users, the negotiation results were different in the two contexts (i.e., a different agreement was reached). These different results highlight how the outcome of the negotiation between the same pair of users may be influenced by the context where the negotiation happens since, as explained in Sections 3 and 4, the grades in a given user’s ethical profile for different contexts may differ. For instance, by referring to Alice and Bob’s personas reported in Figure 5, in the airport context, no agreement was reached as none of them demonstrates a disposition towards giving precedence to the other, according to the personas description, the profiles built upon it, and the users statuses in that context. Instead, in the hospital context, Alice won the negotiation due to the Bob’s disposition in giving precedence to injured people or with medical needs.
@@ 7.2 EQ3: Scalability experiments
To thoroughly assess the scalability of the negotiation process in the scope of EQ3, we performed experiments to measure its overhead in more complex scenarios aptly designed to stress the negotiation process far beyond real-world scenarios. We randomly generated multiple sets of users with profiles consisting of an increasing number of $n$ graded dispositions and an increasing number of $p$ active conditions in their statuses, each of them activating exactly one disposition. We generated 10 different profiles for each value of $n$ in $\{$25, 50, 100, 200, 300}\} and 10 different statuses randomly selecting the active conditions set to true, starting with $p$ equal to 10% of $n$, up to 25%, 50%, 75%, and 100%. This allowed us to experiment with negotiations involving sets of offers of different sizes, both in the number of offers in the set $\mathcal{O}$, and in the number of status conditions $S^{\Join}_{i}$ disclosed with each offer $o_{i}\in\mathcal{O}$. To account for the implicit variance in the randomly generated inputs, we ran the negotiations among the 100 possible combinations of the 10 random users (also allowing two equal users to negotiate) for each combination of $n$ and $p$. As explained above, we ran the negotiation on the two different on-robot and off-robot settings.
@@
It is worth discussing that, in general, the settings where the negotiation time introduced a high overhead, which can be arguably considered unacceptable (e.g., more than 3 seconds) are arguably unlikely in a real-world scenario. As explained above, we ran these settings to stress the system beyond reasonable human-provided input sizes with the primary aim of assessing the scalability dimension. In fact, the aforementioned settings feature huge profiles that consider more than 100 dispositions and, more importantly, an extremely wide and detailed set of possible user conditions (100 or more), most of which hold (i.e., most of the user status conditions are set to true) in the current negotiation context (75% or more). From a practical point of view, this means that we are admitting the existence of users who activated, at the same time, most of the (hundreds) possible status conditions, even though some of them may be conflicting, e.g., child and elderly. As a reference, to make it realistic, personas described in Section 7.1 have, at most, 8 active conditions out of the 27 defined ($\sim$30% of the total). Overall, among the 1500 negotiation runs in the on-robot deployment, 76.5% of them ended in less than 3 seconds; this value raises to 84.73% in the off-robot deployment over the same configurations (among the full set of 2500 negotiation runs in the off-robot, 62.3% of them ended within 3 seconds). However, the results discussed above hint at possible strategies to reduce the negotiation time even in the “extreme” settings where the negotiation overhead becomes unacceptable. Due to the nature of the overhead, which is strongly dependent on the network communication, the off-robot deployment allows reducing the negotiation time if compared to the on-robot deployment (see Table 2). In our experimental setting, the overhead is approximately halved when the negotiation is run locally on an external computer. Moreover, reducing the size of the offer set or the number of exchanged offers would be effective in further reducing the overhead. This can be achieved either by (i) disclosing more than one status condition in each offer or negotiation round, e.g., increasing the value of $i$ in Algorithm 1 and 1 by more than 1 each round, or (ii) introducing a time deadline, as further discussed in Section 9.2.
@@ 8 Discussion
In this section, we discuss the more relevant aspects of the proposed approach.